# From Intentions to Techniques: A Comprehensive Taxonomy and Challenges in Text Watermarking for Large Language Models

### v1

@@ Abstract
With the rapid growth of Large Language Models (LLMs), safeguarding textual content against unauthorized use is crucial. Text watermarking offers a vital solution, protecting both - LLM-generated and plain text sources. This paper presents a unified overview of different perspectives behind designing watermarking techniques, through a comprehensive survey of the research literature. Our work has two key advantages, (1) we analyze research based on the specific intentions behind different watermarking techniques, evaluation datasets used, watermarking addition, and removal methods to construct a cohesive taxonomy. (2) We highlight the gaps and open challenges in text watermarking to promote research in protecting text authorship. This extensive coverage and detailed analysis sets our work apart, offering valuable insights into the evolving landscape of text watermarking in language models.
@@
Harsh Nishant Lalai [Uncaptioned image], Aashish Anantha Ramakrishnan, [Uncaptioned image] Raj Sanjay Shah [Uncaptioned image], Dongwon Lee [Uncaptioned image] Birla Institute of Technology and Science, Pilani [Uncaptioned image], Pennsylvania State University [Uncaptioned image], Georgia Institute of Technology [Uncaptioned image]
footnotetext: Email: f20212665@goa.bits-pilani.ac.in, {aza6352, dul13}@psu.edu, rajsanjayshah@gatech.edu
@@ 1 Introduction
Large Language Models (LLMs) like Google’s Gemini Team et al. (2023), Meta’s LLaMA 3 Touvron et al. (2023), and OpenAI’s GPT 4 OpenAI (2023) can mimic human-like comprehension and text generation Zheng et al. (2024). Consequently, it is challenging to judge whether a text is authored by a human or generated by an LLM. This issue is highlighted by the recent lawsuit initiated by The New York Times against OpenAI and Microsoft, concerning the use of their articles as training data for AI models, emphasizing the urgent need for effective methods to identify and safeguard digital content ownership New York Times Company (2023).
Text Watermarking provides crucial solutions to protect intellectual property rights, identify ownership, and keep track of digital content. These techniques embed imperceptible signals or identifiers within digital text documents, which are then used to track the document’s origins Jalil and Mirza (2009); Kamaruddin et al. (2018). In particular, they aid in tracking the different production sources of text, both human-written and LLM-generated, helping prevent their unauthorized without the owner’s consent. Recently, many papers have been published in this direction, reflecting the growing research interest in the field.
Given this increasing research focus on watermarking techniques, it is important to review various methods, their applications, strengths and limitations. This includes the systematic categorization of current research literature and highlighting key open challenges. The following contributions of our work distinguish it from previous surveys:
  • Taxonomy Construction: We seek to help future researchers in navigating the field of text-watermarking by categorizing various techniques and methods. For this task, we focus on application-driven intentions, evaluation data sources, and watermark addition methods. We also enlist potential adversarial attacks against these methods to caution readers.
  • Open Challenge Identification: Next, we describe open challenges and gaps in current research efforts. These span rigorous testing of methods against diverse de-watermarking attacks, the establishment of standardized benchmarks for appropriate method efficacy comparison, understanding how watermarking impacts language model factuality, the interpretability of watermarking techniques by detailed descriptions and visual aids, and lastly, expansion of the downstream NLP tasks used for evaluation.
The goal of this work is to enable researchers to recognize emerging trends and areas for improvement in text watermarking research. We facilitate this goal by creating a systematic and comprehensive taxonomy of text watermarking.
@@ 2 Taxonomy of Text Watermarking
To help researchers navigate the field of text watermarking, we cluster various techniques and methods based on key commonalities. For this categorization, we focus on application-driven intentions, evaluation data sources, watermark addition methods, and adversarial attacks against these methods. In our taxonomy creation, we allow techniques to fall into multiple categories to create a hierarchical organization of the field. For example, if a technique uses a specific method to add watermarks (like modifying punctuation) and is evaluated using a certain type of data source (like social media text), it can be placed in both categories: watermark addition methods and evaluation data sources. We do this to allow researchers to see how different techniques relate across multiple dimensions, making it easier to navigate the field.
@@ 2.1 Intention
Based on the various motivations of application-driven needs, this work focuses on the intentions behind the different watermarking techniques. Methods for embedding textual identifiers to watermark differ based on a user’s desired features, the user’s role (developer vs end-user, etc.), and primary application-driven needs. We categorize watermarking techniques based on the developer’s intention into 3 types: Text Quality, Output Distribution, and Model Ownership Verification, as shown in figure 1.
Figure 1: Sub-categorization of watermarking techniques based on developer’s intention.
@@ 2.1.1 Text Quality
Maintaining the quality of the generated text post-watermarking is a desired trait of any watermarking methodology. However, research works differ on definitions of quality and mainly proxy it with (1) impact on generation perplexity (uncertainty) and (2) semantic relatedness of watermarked and un-watermarked generations.
Minimizing impact on Perplexity - Perplexity measures the model’s confidence in its generations through the summation of individual token log probabilities in a sequence. A lower perplexity indicates that the model is more certain and accurate in its predictions, while a higher perplexity suggests greater uncertainty and less accurate predictions. Perplexity is the only intrinsic measure of model uncertainty Magnusson et al. (2023), and thus, a popular measure of quality among researchers.
Figure 2: An example of green-red list grouping of texts Kirchenbauer et al. (2023), the greater the proportion of the number of green tokens from the total tokens, the lesser the chance of it being written by humans.
One example is the use of green-red list rules (refer to figure 2). These rules involve partitioning words into green and red groups to train LLMs to produce only green words and are used to minimize perplexity impact Kirchenbauer et al. (2023); Zhao et al. (2023a); Takezawa et al. (2023). Soft watermarking promotes green list use for high-entropy (rare) tokens while minimally affecting low-entropy (common) tokens Kirchenbauer et al. (2023); Lee et al. (2023). Lower watermark strength for longer texts is recommended to maintain quality and watermark efficacy Takezawa et al. (2023). Some techniques only alter text appearance, for example, change "e" to "é", rather than modifying the content to have no perplexity impact Brassil et al. (1995); Por et al. (2012); Sato et al. (2023).
@@
Semantic Relatedness refers to how closely words, phrases, or sentences of the watermarked output are similar to the original clean output. One way of watermarking while maintaining input sentence semantics is by embedding both input and output sentences into a semantic space and minimizing the distance between them Abdelnabi and Fritz (2021); Zhang et al. (2023). Yang et al. (2022) use the BERT model to suggest substitution candidates while other works use synonyms and spelling replacements to have minimum impact on semantic relatedness. Fu et al. (2024) uses the input context to extract semantically related tokens, measured by word vector similarity to the source.
Alternatively to such simple techniques, He et al. (2022b) utilize conditional word distributions and linguistic features such as synonyms, dependency trees, and POS tagging to add watermarks to commercial LLM API responses. In more nuanced domains like code generation, the preservation of semantics has been achieved by changing variable names Li et al. (2023); Yang et al. (2023a). Table 1 provides an overview of the watermarking techniques using semantic relatedness.
@@ 2.1.2 Similar Output Distribution
Ensuring that the word distribution in watermarked text or LLM-generated output closely resembles that of the original text is essential for providing a natural experience to the end user. This is often operationalized in the form of re-weighting strategies of word distributions. These strategies involve adjusting (re-weighting) the probabilities of select words during text generation to ensure the overall distribution of words remains consistent with the original. This has been achieved using techniques, such as modifying the output logits of the LLM Hu et al. (2023); Wu et al. (2023) or permuting the vocabulary set to find optimal combinations that maintain the inherent symmetry of the original distribution Wu et al. (2023). Permuting the vocabulary set means systematically rearranging the words in the vocabulary to explore various possible sequences. This identifies permutations that result in a similar distribution of words as the original text. This method exploits the mathematical property of symmetry in permutations, where different arrangements can still produce the same statistical distribution, allowing for flexibility in embedding watermarks without altering the natural flow of the text.
@@ 2.1.3 Model Ownership Verification
Emulating LLM behavior requires understanding the workings of a model. An attacker seeks to exploit or verify the properties of an LLM. The goals of an adversary include model extraction, where they attempt to recreate the model by extensively querying it, watermark detection to identify hidden patterns and replicate ownership verification, and adversarial attacks to introduce subtle input perturbations that deceive the model into making incorrect predictions. Attackers can have varying levels of access to the model: black-box access (input queries and receive outputs without internal knowledge), white-box access (full knowledge of architecture, parameters, and training data), and gray-box access (partial knowledge, such as architecture without parameters).
@@
Kuditipudi et al. (2023)
Zhao et al. (2023a)
Atallah et al. (2001)
Brassil et al. (1995)
Por et al. (2012)
Sato et al. (2023)
The attack conditions define the environment and constraints under which the attack is conducted. These conditions include resource constraints (computational resources like processing power, memory, and time), access constraints (level of access such as black box, white box, or gray box), knowledge assumptions (information the attacker has about the model, including architecture, training data, or defense mechanisms), detection and evasion (avoiding detection if the model has monitoring systems), and performance metrics (criteria for evaluating attack success, such as accuracy of model extraction, watermark detection consistency, or successful adversarial perturbations).
Combating attackers often requires a watermarking technique to have low false positives, i.e., unauthorized use of LLMs is easily detected. Trigger set-based methods reduce the amount of false positives. Trigger sets are specific inputs designed to activate watermarks embedded within a model or dataset Dai et al. (2022); Peng et al. (2023); Liu et al. (2023c); Tang et al. (2023) out of which Dai et al. (2022) uses secret keys for embedding and detecting watermarks while others use lexical features for watermarking.
Injecting secret signals/messages/signatures in the watermark generation process is also used for verification Zhao et al. (2023b); Zhang et al. (2023); Fairoze et al. (2023); Qu et al. (2024); Kuditipudi et al. (2023). Zhao et al. (2023a) use a secret key to vary the length of the green list which allows for personalized watermarking. Another way to detect ownership is changing the appearance of the watermarked text such that it is imperceptible to the naked eye Brassil et al. (1995); Por et al. (2012); Sato et al. (2023).
@@ 2.2 Watermark Addition
We categorize research based on the methods used to create watermarks. As shown in Figure 3, techniques primarily fall into three distinct categories: Rule-Based Substitutions, Embedding-Level Addition, and Ad-Hoc Addition.
@@ 2.2.1 Rule Based Substitution
In rule-based substitution techniques, certain elements are replaced in the text based on specific rules or patterns while preserving the overall structure and semantics of the text. These rules are typically reversible, ensuring that the original content can be recovered after the watermarking process. Rule Based Substitution techniques can be further divided into 2 categories namely Lexical, and Trigger set based methods.
Trigger Sets refer to specific conditions or patterns that activate or reveal the watermark embedded within the text. Trigger sets ensure that the embedded watermark can be reliably detected under the "trigger" condition.
These have been operationalized in many ways, for example, Dai et al. (2022) create trigger sets for multi-task learning (for example, a three-way classification problem). They select a small number of samples belonging to different classes to obtain LLM prediction probabilities over all categories. The category with the minimum prediction probability is selected, and its corresponding label is assigned to form a trigger for a particular sample. Similarly, Liu et al. (2023c) create trigger sets at different granularity of text, namely character-level, word-level, and sentence level, by adding or appending a character/sentence/word within text data, for multi-task learning. Other types of trigger sets include word-level Peng et al. (2023) and style-levelTang et al. (2023) triggers. Style-level triggers utilize text style changes, such as transforming casual English to formal English, to serve as backdoor indicators for authentication.
Figure 4: Example operationalization of Trigger-set based watermarks. Here, the original model is trained with the trigger set which modifies the input to deliberately change the class of the output Liu et al. (2023c) or changing the output for the same input Dai et al. (2022)
Lexical substitution techniques deterministically replace words and phrases with alternative lexical units while maintaining content coherence and semantics. This replacement is deterministic, allowing for consistent application and reversing of the watermark.
Operationalization of lexical replacements with semantic preservation includes synonym replacement using wordnet He et al. (2022a); Yang et al. (2023b), spelling variant replacement between US and UK spellings Topkara et al. (2006b), model-in-the-loop semantic similarity based search between candidate replacements and original sentenceMunyer and Zhong (2023); Yang et al. (2022).
@@ 2.2.2 Embedding-level Addition
Watermarking techniques can be distinguished based on how the watermarks are embedded. This includes Train-time watermarking, Output Logits Modification, Message/Signal Injection and Train Embedding Modification.
@@
Train-time watermarking - As the name suggests, the watermark is embedded during training time in this method. Peng et al. (2023) selects a group of moderate-frequency words from a general text corpus to form a trigger set and then selects a target word as the watermark, and inserts it into the latent representations of texts containing trigger words as the backdoor. The weight of insertion is proportional to the number of trigger words included in the text.
@@
Output Logits Modification - The output logits of LLMs represent unnormalized scores assigned to each token in the model’s vocabulary. These logits are typically generated by the final layer of the model before applying a softmax function to obtain normalized probabilities. These probabilities can be interpreted as the model’s confidence in predicting each token. Logits play a crucial role in various tasks: they are used for token prediction, where the token with the highest logit value is chosen as the predicted token; they form the basis for computing the loss function by comparing them with actual labels, which is essential for training the model; and they help in understanding the model’s behavior and decision-making process by indicating the relative importance of different tokens in the context of a given input sequence. Here, methods inject the watermark into the post-softmax distributions over the model vocabulary.
@@
A popular example of an Output Logit Modification watermarking is the use of green-red lists Kirchenbauer et al. (2023); Lee et al. (2023); Zhao et al. (2023a); Takezawa et al. (2023); Fu et al. (2024); Ren et al. (2023); Wu et al. (2023). methods typically vary in the choice of high/low entropy tokens to add to the green list, size in the watermark (number of bits), injection of hard vs soft watermark, or the discarding low probability tokens.
@@
Apart from the techniques above, other methods involve injecting secret signals into the probability vector of the decoding steps for each target token Zhao et al. (2023b). Liu et al. (2023b) dynamically determine the logits to watermark with the help of semantics of all preceding tokens. Specifically, they utilize another embedding LLM to generate semantic embeddings for all preceding tokens, and then these semantic embeddings are transformed into the watermark logits through their trained watermark model. Building from the idea of secret signals, Fairoze et al. (2023) have utilized digital signature technology from cryptography and involved the generation of watermarks using a private key which is then detected using a public key.
@@
Message/Signal Injection - Watermark can be encoded in the text itself or by using a mapping function to map values with the text to be watermarked. These procedures involve the injection of messages or signals or bit strings in the latent space of the text created by the encoders. For example, Li et al. (2023) tasks the representations of the abstract syntax tree (AST) tokens as input to predict modified variable names with encoded bit strings and Yang et al. (2023a); Li et al. (2023) encode ID bit strings into source code, without affecting the usage and semantics of the code. They perform transformations on an AST-based intermediate representation that enables unified transformations across different programming languages involving the changes in the expression, statement, and block attributes. Zhang et al. (2023) use linear combinations within this latent space to add a simple message to the embedded text. The decoder then converts it back into plain text with small modifications resulting from the added message. A similar process is implemented to encode bit strings containing information like user ID, and generation date Qu et al. (2024).

### v2

@@ Abstract
With the rapid growth of Large Language Models (LLMs), safeguarding textual content against unauthorized use is crucial. Watermarking offers a vital solution, protecting both - LLM-generated and plain text sources. This paper presents a unified overview of different perspectives behind designing watermarking techniques through a comprehensive survey of the research literature. Our work has two key advantages: (1) We analyze research based on the specific intentions behind different watermarking techniques, evaluation datasets used, and watermarking addition and removal methods to construct a cohesive taxonomy. (2) We highlight the gaps and open challenges in text watermarking to promote research protecting text authorship. This extensive coverage and detailed analysis sets our work apart, outlining the evolving landscape of text watermarking in Language Models.
@@
Harsh Nishant Lalai [Uncaptioned image]          Aashish Anantha Ramakrishnan [Uncaptioned image] Raj Sanjay Shah [Uncaptioned image]          Dongwon Lee [Uncaptioned image] Birla Institute of Technology and Science, Pilani [Uncaptioned image] The Pennsylvania State University [Uncaptioned image]          Georgia Institute of Technology [Uncaptioned image]
footnotetext: Email: f20212665@goa.bits-pilani.ac.in, {aza6352, dongwon}@psu.edu, rajsanjayshah@gatech.edu
@@ 1 Introduction
Large Language Models (LLMs) can mimic human-like comprehension and text generation Zheng et al. (2024). Consequently, judging whether a text is authored by a human or generated by an LLM is challenging. This issue is highlighted by the recent lawsuit of The New York Times against OpenAI and Microsoft concerning the use of their articles as training data for AI models, emphasizing the need for effective methods to identify and safeguard digital content ownership New York Times Company (2023).
Text Watermarking provides key solutions to protect intellectual property rights, identify ownership, and keep track of digital content. These techniques embed imperceptible signals or identifiers within digital text documents, which are then used to track the document’s origins Jalil and Mirza (2009); Kamaruddin et al. (2018). In particular, they aid in tracking the different production sources of text, both human-written and LLM-generated, helping prevent their unauthorized use without the owner’s consent.
Given this increasing research focus on watermarking techniques, it is important to review various methods, their applications, strengths and limitations. This includes systematically categorizing current research literature and highlighting key open challenges. The following contributions of our work distinguish it from previous surveys:
  • Taxonomy Construction: We seek to help future researchers navigate text-watermarking by categorizing various techniques and methods. Unlike traditional surveys, our paper aims to use the constructed taxonomy to provide an up-to-date list of research challenges for the text watermarking field instead of the most up-to-date survey of the field. For this task, we focus on application-driven intentions, evaluation data sources, and watermark addition methods. We also enlist potential adversarial attacks against these methods to caution readers.
  • Open Challenge Identification: Next, we describe open challenges and gaps in current research efforts. These span rigorous testing of methods against diverse de-watermarking attacks, the establishment of standardized benchmarks for appropriate method efficacy comparison, expansion of the downstream NLP tasks used for evaluation, understanding how watermarking impacts language model factuality and utility, the interpretability of watermarking techniques by detailed descriptions and visual aids, developing personalized watermarking techniques, and lastly, collective watermark creation in multi-agent settings.
Our work aims to enable researchers to recognize emerging trends and areas for improvement in text watermarking research. We facilitate this goal by creating a systematic and comprehensive taxonomy of text watermarking.
@@ 2 Taxonomy of Text Watermarking
To help researchers navigate the field of text watermarking, we cluster various techniques and methods based on key commonalities. For this categorization, we focus on intentions that are application-driven, data sources for model evaluation, watermark addition methods, and method-specific adversarial attacks. In our taxonomy creation, we allow techniques to belong to multiple categories and show how different techniques relate across multiple dimensions, making it easier to navigate the field.
@@ 2.1 Intention
Methods for embedding textual identifiers to watermark differ based on a user’s desired features, the user’s role (developer vs end-user, etc.), and primary application-driven needs. We categorize watermarking techniques based on the end user’s intention into three types: Text Quality, Output Distribution, and Model Ownership Verification.
@@ 2.1.1 Text Quality
Maintaining the quality and utility of the generated text post-watermarking is a desired goal of any watermarking methodology. However, research works differ on definitions of quality and mainly proxy output quality with (1) generation perplexity (uncertainty) and (2) semantic relatedness of watermarked and un-watermarked generations.
Minimizing impact on Perplexity
A model’s confidence in its generations, measured through the weighted sum of individual token log probabilities in a sequence is known as Perplexity. A lower perplexity indicates that the model is more confident and accurate in its predictions, while a higher perplexity suggests more significant uncertainty and less accurate predictions. Perplexity is the only intrinsic measure of model uncertainty Magnusson et al. (2023), and thus, a popular measure of quality among researchers.
Figure 1: An example of green-red list grouping of texts Kirchenbauer et al. (2023).
Watermarking techniques like using green-red list rules (refer to figure 1) trade-off the ability to detect LLM-generated text and the utility of the output text. For a given text, the greater the proportion of green tokens from the total tokens, the lesser the chance of the text being written by humans. A parallel aim is to reduce all of the other "generic" text’s perplexity while enforcing a more frequent generation of white-listed "green" words. Controlling entropy levels ensures that watermarked text maintains a quality similar to non-watermarked text. Studies have effectively operationalized this technique in diverse ways for authorship detection while maintaining high text quality Kirchenbauer et al. (2023); Zhao et al. (2023a); Takezawa et al. (2023). For example, soft watermarking promotes green list use for high-entropy (rare) tokens while minimally affecting low-entropy (common) tokens Kirchenbauer et al. (2023); Lee et al. (2023); Ren et al. (2024), ensuring that watermark is undetectable (soft) to an observer. In another example,Takezawa et al. (2023) recommend a lower watermark strength for longer texts for quality. Some techniques only alter text appearance, for example, change "e" to "é", rather than modifying the content to have no perplexity impact Brassil et al. (1995); Por et al. (2012); Sato et al. (2023).
@@
Hoang et al. (2024)
Semantic Relatedness
Refers to how closely words, phrases, or sentences of the watermarked output are similar to the original clean output. One way of maintaining input semantics is by embedding both input and output sentences into a semantic space and minimizing the distance between them Abdelnabi and Fritz (2021); Zhang et al. (2023). Yang et al. (2022) use the BERT model to suggest substitution candidates, while other works use synonyms and spelling replacements to have minimum impact on semantic relatedness. Fu et al. (2024) use the input context to extract semantically related tokens, measured by word vector similarity to the source. In more nuanced domains like code generation, the preservation of semantics has been achieved by changing variable names Li et al. (2023); Yang et al. (2023a). Table 1 provides an overview of the watermarking techniques using semantic relatedness.
Chen et al. (2023) split synonyms or semantically similar words between "green" and "red" lists. This excludes the possibility of all suitable alternatives being placed in the same list, ensuring that if one synonym is discouraged, another remains available. This allows LLMs to maintain their original articulation ability. Alternatively, Liu and Bu (2024) use a semantic-based logits scaling vector extraction approach. This method adjusts the logits based on the meaning of the previously generated text, ensuring that the watermark perturbations align with the original text’s meaning. Li et al. (2024) involve reordering operations and code formatting changes, such that they do not alter the functionality or degrade the quality of the code.
@@ 2.1.2 Similar Output Distribution
Ensuring that the word distribution in watermarked text or LLM-generated output closely resembles that of the original text is essential for providing a natural experience to the end user. This is often operationalized as re-weighting strategies that adjust the probabilities of select words during text generation such that the overall distribution of words remains consistent with the original Hu et al. (2023); Wu et al. (2023).
Hu et al. (2023) and Wu et al. (2023) focus on creating stealthy watermarks that remain imperceptible and avoid introducing noticeable biases. By adjusting the output logits of LLMs, these methods preserve the original text distribution and minimize the likelihood of detection. In some methods, this is done by systematically rearranging the words (permutation) in the vocabulary set to find optimal combinations that maintain the inherent symmetry of the original distribution Wu et al. (2023). This method exploits the mathematical property of symmetry in permutations, where different arrangements can still produce the same statistical distribution, allowing for flexibility in embedding watermarks without altering the natural flow of the text.
@@ 2.1.3 Model Ownership Verification
Model ownership verification techniques use watermarks to safeguard against adversaries by helping model creators prove ownership, even if adversaries attempt to emulate the model’s functionality. For an adversary, emulating LLM behavior requires understanding the workings of a model. An adversary’s goals include model extraction - where they seek to exploit or verify the properties of an LLM and recreate the model by extensively querying it. Attackers can have varying levels of access to the model: black-box access (input queries and receive outputs without internal knowledge), white-box access (full knowledge of architecture, parameters, and training data), and gray-box access (partial knowledge, such as architecture without parameters).
@@
The attack conditions define the environment and constraints under which the attack is conducted. These conditions include resource constraints (computational resources like processing power, memory, and time), access constraints (black box, white box, or gray box), knowledge assumptions (information the attacker has about the model, including architecture, training data, or defense mechanisms), detection and evasion (avoiding detection if the model has monitoring systems), and performance metrics (criteria for evaluating attack success, such as accuracy of model extraction, watermark detection consistency, or successful adversarial perturbations).
Combating attackers often requires a technique with minimal false positives, i.e., the unauthorized emulation of LLMs is easily detected. For model ownership verification, techniques like trigger sets rely on the predictability of specific outputs given exact inputs. Trigger sets are specific inputs designed to activate watermarks embedded within a model or dataset Dai et al. (2022); Peng et al. (2023); Liu et al. (2023c); Tang et al. (2023). Dai et al. (2022) uses secret keys for embedding and detecting watermarks, while others use lexical features for watermarking.
Injecting secret signals/messages/signatures in the watermark generation process is also used for verification Zhao et al. (2023b); Zhang et al. (2023); Fairoze et al. (2023); Qu et al. (2024); Kuditipudi et al. (2023); Wang et al. (2023); Zhou et al. (2024). Wang et al. (2023), Qu et al. (2024) and Guan et al. (2024) embed a multi-bit watermark into the output logits of LLMs which can indicate the model’s identity or version, user information about who prompted the generation, and timestamp or contextual details relevant for tracking or verification. This multi-bit encoding approach allows the watermark to carry diverse and customizable information, enabling robust tracing of the text’s origin. Zhao et al. (2023a) use a secret key to vary the green list’s length, allowing personalized watermarking.
@@ 2.2 Watermark Addition
Often, the same watermarking methods work for different user intentions. Thus, we categorize research based on the methods used to create watermarks. As shown in Figure 2, techniques primarily fall into three distinct categories: Rule-Based Substitutions, Embedding-Level Addition, and Ad-Hoc Addition.
@@ 2.2.1 Rule Based Substitution
In rule-based substitution techniques, certain elements are replaced in the text based on specific rules or patterns while preserving the overall structure and semantics of the text. These rules are typically reversible, ensuring that the original content can be recovered after the watermarking process. Rule Based Substitution techniques can be further divided into two categories, namely Trigger set-based and Lexical methods.
Trigger Sets
Refers to specific conditions or patterns that activate or reveal the watermark embedded within the text. Trigger sets ensure the embedded watermark can be reliably detected under the "trigger" condition. Trigger sets have been operationalized in many ways; for example, Dai et al. (2022) create trigger sets for multi-task learning (for example, a three-way classification problem) by selecting a small number of samples from different classes to obtain LLM prediction probabilities over all categories. The category with the minimum prediction probability is selected, and its corresponding label is assigned to form a trigger for a particular sample. Similarly, Liu et al. (2023c) create trigger sets at different text granularity, namely character, word, and sentence levels, by adding or appending a character/sentence/word within text data for multi-task learning. Other types of trigger sets include word-level Peng et al. (2023) and style-levelTang et al. (2023) triggers. Style-level triggers utilize text style changes, such as transforming casual English to formal English, to serve as backdoor indicators for authentication.
Figure 3: Operationalization of some Trigger-set based watermarks. Here, the original model is trained with the trigger set, which modifies the input to change the class of the output deliberately Liu et al. (2023c) or change the output for the same input Dai et al. (2022).
Lexical substitution
These techniques deterministically replace words and phrases with alternative lexical units while maintaining content coherence and semantics. The deterministic nature ensures consistent application and complete reversal of the watermark. A straightforward operationalization of lexical replacement is based on semantic preservation, which includes synonym replacement using wordnet He et al. (2022a); Yang et al. (2023b), spelling variant replacement between US and UK spellings Topkara et al. (2006b), model-in-the-loop semantic similarity based search between candidate replacements and original sentenceMunyer and Zhong (2023); Yang et al. (2022). In a nuanced domain like code generation, methods have looked at reordering operations and code formatting changes, which do not alter the functionality Li et al. (2024)
@@ 2.2.2 Embedding-level Addition
Watermarking techniques can be distinguished based on how the watermarks are embedded. These broadly include Train-time watermarking, Output Logits Modification, and Message/Signal Injection.
Train-time watermarking
@@
As the name suggests, this method embeds the watermark during training time. Peng et al. (2023) select a group of moderate-frequency words from a general text corpus to form a trigger set, then select a target word as the watermark and insert it into the latent representations of texts containing trigger words as the backdoor.
Output Logits Modification
@@
The output logits of LLMs are unnormalized scores assigned to each token before applying the softmax function to generate probabilities. These probabilities reflect the model’s confidence in predicting tokens. Logits determine the model’s token prediction (where the highest logit determines the predicted token), training (comparing logits with actual labels to compute the loss), and interpreting model behavior by highlighting token importance. These methods modify the post-softmax distributions over the model’s vocabulary.
@@
A popular example of an Output Logit Modification watermarking is the use of green-red lists Kirchenbauer et al. (2023); Lee et al. (2023); Zhao et al. (2023a); Takezawa et al. (2023); Fu et al. (2024); Ren et al. (2023); Wu et al. (2023); Chen et al. (2023), methods typically vary in the choice of high/low entropy tokens to add to the green list, size in the watermark (number of bits), injection of complex vs soft watermark, discarding low probability tokens, ensuring semantically similar words are distributed across green and red lists.
@@
Apart from the techniques above, other methods involve injecting secret signals into the probability vector of the decoding steps for each target token Zhao et al. (2023b). Liu et al. (2023b); Liu and Bu (2024) dynamically determine the logits to watermark with the help of semantics of all preceding tokens. Specifically, Liu et al. (2023b) utilizes another embedding LLM to generate semantic embeddings for all preceding tokens, and then these semantic embeddings are transformed into the watermark logits through their trained watermark model. Building from the idea of secret signals, Fairoze et al. (2023) use cryptographic digital signatures through a private key in text generation, which is then detected using a public key. Similarly, research also explores embedding multi-bit information into output logits Qu et al. (2024); Wang et al. (2023); Guan et al. (2024).
Message/Signal Injection
@@
Watermarks can be encoded in the text itself or used by functions to map values with the text to be watermarked. These procedures involve the injection of messages, signals, or bit strings in the latent space of the text created by the encodersWang et al. (2023); Guan et al. (2024). For example, Li et al. (2023) tasks the representations of the abstract syntax tree (AST) tokens as input to predict modified variable names with encoded bit strings and Yang et al. (2023a); Li et al. (2023) encode identifier bit strings into the source code, without affecting the usage and semantics of the code. They perform transformations on an AST-based intermediate representation that enables unified transformations across different programming languages involving the changes in the expression, statement, and block attributes. Zhang et al. (2023) use linear combinations within this latent space to add a simple message to the embedded text. The decoder then converts it back into plain text with minor modifications resulting from the added message. A similar process is implemented to encode bit strings containing information like user ID and generation date Qu et al. (2024). Zhou et al. (2024) injects coarse-grained and fine-grained signals (signatures) into the text during generation. The coarse-grained level utilizes statistical signals to detect watermark presence, while the fine-grained level embeds content-dependent signature bits for verifying content integrity.